# Evaluating The Explainability of State-of-the-Art Machine Learning-based Online Network Intrusion Detection Systems

### v2

@@ Abstract
Network Intrusion Detection Systems (NIDSs) which use machine learning (ML) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying ML-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art ML-based online NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP). Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some ML-based NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to inductive bias than other models.
@@ I Introduction
Securing networks against intrusion is paramount, and network intrusion detection systems (NIDS) play a pivotal role in achieving this. Traditionally, NIDSs have relied heavily on signature-based methods [5, 6]- essentially matching attack patterns to known signatures. However, these approaches demand significant manual effort from experts to create and maintain those signatures. Moreover, they fall short when it comes to identifying novel, previously unseen attacks. With the introduction of machine learning (ML)-based NIDSs which leverage ML models to enhance existing defenses, impressive performance and accuracy has been achieved [7, 8]. Unlike their signature-centric counterparts, ML-based NIDSs do not require experts to painstakingly prepare signatures in advance. In fact, unsupervised ML-based NIDSs can even raise alerts for hitherto unknown attacks. Thanks to these advantages and the rapid evolution of ML techniques, ML-based NIDSs have made significant strides in recent years.
@@
Despite the promise of ML-based NIDSs, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying them in real-world production environments [9]. The crux of the issue lies in the black-box nature of many of these proposed solutions. Essentially, their inner workings (how and why they make decisions) remain elusive, unlike simpler but less effective rule-based approaches that security professionals are accustomed to. Security analysts rely on ML-based NIDSs for critical decisions. However, without clear explanations for the decisions made by these underlying ML models, it becomes difficult for the analysts to trust them. Moreover, ML often operates as a black box within the NIDS. Consequently, hidden correlations stemming from unrelated artifacts can slip past undetected, leading to a flawed perception of the system’s capabilities. Explanation tools and techniques can unveil those hidden correlations, allowing practitioners to assess their impact on the NIDS’s overall effectiveness. As it stands, the current state-of-the-art online NIDSs lack inherent explainability. Therefore, there is a pressing need to scrutinize them using these very explanation techniques.
@@
In this work, we analyze state-of-the-art ML-based online NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [10], SHAP [11]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Using both global explanations and local explanations for the models’ decisions, the most prominent features used by each ML-based NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our devised qualitative approach) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art ML-based online NIDSs other than Kitsune using xAI techniques.
@@
  • We analyze state-of-the-art ML-based online NIDS models using explainable AI techniques.
  • We present the most prominent features used by each NIDS model considered to make decisions on input data samples using the explanations generated earlier.
  • We compare the explanations generated across the models for a given xAI method as well as the explanations generated across xAI methods for a given model.
  • We evaluate the vulnerability of each NIDS model to inductive bias.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on ML-based security systems and explainable ML-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art ML-based online NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the ML-based NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the online NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
@@ II Related Work
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [14, 15, 16, 17] have employed xAI methods such as LIME [18] and SHAP [11] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to network intrusion datasets (NSL-KDD dataset [22], CICIDS-2017 [23]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [24], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage deep learning techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
@@
A few works have also proposed ML-based attack detection systems with built-in explainability. In [27], the authors have implemented an xAI solution to provide accurate detection and classification of the DNS-over-HTTPS attacks. It is based on a balanced stacked random forest classifier. The authors have highlighted the underlying feature contributions to provide transparent and explainable results from the model. In [28], the authors have proposed an NIDS for IT networks combining ensemble learning and stacking with a meta-learner (CNN) that works on graphical representation of traffic flows and provides the required explainability level for the decisions made. They also provide visual representations of network anomalies that allows security analysts to interpret and gain insights into the detected network anomalies.
Our work differs from the previous literature in the following aspects:
  • Existing works that evaluate xAI methods on ML-based NIDSs do not focus on state-of-the-art online NIDSs and do not include the latest xAI tools for ML-based network security such as TRUSTEE ([14, 15, 16, 17]).
  • Existing explainable-by-design AI-based attack/intrusion detection systems are targeted at specific attacks (DNS-over-HTTPS attacks in [27]).
@@ III Background
In this paper, since we delve into explaining the decisions of state-of-the-art ML-based online NIDSs, it is essential to briefly explore the types of explanation methods present in academic literature, the NIDSs that we consider for explanation in this work and the various aspects of our investigation with xAI techniques on those NIDSs.
@@ III-A Explanation Methods- Types
All explanation methods for ML models can be broadly categorized into black-box and white-box explanations.
@@ III-B Prominent Black-box XAI Tools
ML-based NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the ML-based NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
Overview of TRUSTEE: TRUSTEE [10] is a state-of-the-art interpretability tool for ML-based security solutions which generates interpretable decision trees (DT) for a given ML model. It can be used to find out inductive biases in black-box models (e.g., shortcut learning, spurious correlations, sensitiveness to out-of-distribution samples). While generating DT, TRUSTEE focuses on the following desirable properties: model-agnosticism (should be applicable to any given black-box model), model fidelity (how accurate the DT is with respect to the original ML model), model stability (the decision rules should correctly describe how the given black-box model makes a significant number of its decisions and should be largely insensitive to the particular data samples that TRUSTEE used in the process of selecting its final DT explanation) and model complexity (decision rules should be readily recognizable by domain experts). The important aspects of generated DT explanation are summarized in a trust report which helps end users determine whether the given black-box model suffers from the problem of underspecification and can be trusted.
@@
III-C State-of-the-art ML-based Online NIDSs
We now present a brief overview of each state-of-the-art ML-based online NIDS that we investigate using xAI techniques in this work.
Kitsune [32] proposes an online unsupervised anomaly detection system based on an ensemble of autoencoders which is lightweight in terms of memory footprint and meant to be deployed on network gateways and routers to detect attacks on the local network. It uses a damped incremental statistics method to extract network traffic features which has the advantage of extracting features from dynamic network traffic at high speed.
HorusEye [33] is a two-stage anomaly detection framework for IoT. In the first stage, preliminary burst-level anomaly detection is implemented on the data plane. The suspicious traffic is then reported to the control plane for further investigation. To reduce the false positive rate, the control plane carries out the second stage, where more thorough anomaly detection is performed over the reported suspicious traffic using flow-level features and an autoencoder.
ENIDrift [34] addresses real-world issues in ML-based NIDS deployment such as concept drift (change in statistical properties of target variable over time), class-imbalanced training data and well-crafted ML attacks (training data contamination, adversarial attacks). It includes three main components: iP2V- an incremental feature extraction method based on Word2Vec, a sub-classifier generation module and an update module.
@@ III-C State-of-the-art ML-based Online NIDSs
IoTA [36] is an intrusion detection system for IoT devices which uses fully packet-level models to profile IoT device traffic patterns. These models are based on automata constructed from raw device traffic traces whose states represent packets with similar context, representative features are packet length and direction and when grouped together based on connectedness, represent device traffic flows. At runtime, ongoing IoT device traffic is matched with the extracted normal traffic profiles towards intrusion detection. However, the authors have not released the source code for IoTA implementation publicly and therefore, we have not included it in our analysis.
IV xAI Analysis of ML-based Online NIDSs
In this work, we analyze four state-of-the-art ML-based online NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible Online NIDS Implementation
The TRUSTEE trust report generating API has two requirements: (1) a trained ML model with a predict() function taking training/test dataset as the only input and outputting predicted labels, and (2) training/test datasets and corresponding predicted labels should be in numpy array/pandas dataframe format. The kernel SHAP API shares the requirements of TRUSTEE API. Except Kitsune, other NIDS source codes do not provide trained ML model or training/test data in a format which can be directly input to TRUSTEE API or kernel SHAP API.
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art ML-based online NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML models such as the NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.

### v3

# Evaluating The Explainability of State-of-the-Art Deep Learning-based Network Intrusion Detection Systems

@@ Abstract
Network Intrusion Detection Systems (NIDSs) which use deep learning (DL) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying DL-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art DL-based NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP) through extensive experiments with two different attack datasets. Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some DL-based NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to inductive bias than other models.
@@ I Introduction
Securing networks against intrusion is paramount, and network intrusion detection systems (NIDS) play a pivotal role in achieving this. Traditionally, NIDSs have relied heavily on signature-based methods [1, 2]- essentially matching attack patterns to known signatures. However, these approaches demand significant manual effort from experts to create and maintain those signatures. Moreover, they fall short when it comes to identifying novel, previously unseen attacks. With the introduction of deep learning (DL)-based NIDSs which leverage DL models to enhance existing defenses, impressive performance and accuracy has been achieved [3, 4] and they have been found to perform better than conventional machine learning (ML) algorithms [5]. In fact, unsupervised DL-based NIDSs can even raise alerts for hitherto unknown attacks. Thanks to these advantages and the rapid evolution of DL techniques, DL-based NIDSs have made significant strides in recent years.
@@
Despite the promise of DL-based NIDSs, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying them in real-world production environments [6]. The crux of the issue lies in the black-box nature of many of these proposed solutions. Essentially, their inner workings (how and why they make decisions) remain elusive, unlike simpler but less effective rule-based approaches that security professionals are accustomed to. Security analysts rely on DL-based NIDSs for critical decisions. However, without clear explanations for the decisions made by these underlying DL models, it becomes difficult for the analysts to trust them. Moreover, DL often operates as a black box within the NIDS. Consequently, hidden correlations stemming from unrelated artifacts can slip past undetected, leading to a flawed perception of the system’s capabilities. Explanation tools and techniques can unveil those hidden correlations, allowing practitioners to assess their impact on the NIDS’s overall effectiveness. As it stands, the current state-of-the-art DL-based NIDSs lack inherent explainability. Therefore, there is a pressing need to scrutinize them using these very explanation techniques.
@@
In this work, we analyze state-of-the-art DL-based NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [7], SHAP [8]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Through extensive experiments with two different attack datasets and using both global explanations and local explanations for the models’ decisions, the most prominent features used by each DL-based NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our proposed metric) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art DL-based NIDSs other than Kitsune using xAI techniques.
@@
  • We analyze state-of-the-art DL-based NIDS models using explainable AI techniques.
  • We present the most prominent features used by each NIDS model to make decisions on input data samples using the explanations generated earlier.
  • Using the same explanations and some evaluation metrics, we illustrate that:
    • Some DL-based NIDS models can be better explained (in terms of interpretable decision trees) than other models.
    • xAI explanations generated using different tools (e.g., TRUSTEE, SHAP) are in conflict with each other for most of the NIDS models considered.
    • Some NIDS models are more vulnerable to inductive bias (in terms of learnt packet rates) than other models.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on DL-based security systems and explainable DL-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art DL-based NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the DL-based NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the DL-based NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
@@ II Related Work
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [11, 12, 13, 14] have employed xAI methods such as LIME [15] and SHAP [8] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to network intrusion datasets (NSL-KDD dataset [16], CICIDS-2017 [17]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [18], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage DL techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
@@
A few works have also proposed ML/DL-based attack detection systems with built-in explainability. In [21], the authors have implemented an xAI solution to provide accurate detection and classification of the DNS-over-HTTPS attacks. It is based on a balanced stacked random forest classifier. The authors have highlighted the underlying feature contributions to provide transparent and explainable results from the model. In [22], the authors have proposed an NIDS for IT networks combining ensemble learning and stacking with a meta-learner (CNN) that works on graphical representation of traffic flows and provides the required explainability level for the decisions made. They also provide visual representations of network anomalies that allows security analysts to interpret and gain insights into the detected network anomalies.
Though not targeted at generating explanations, recent works on adversarial learning in NIDSs aim to generate specific examples/feature vectors which can evade detection. Based on a complete/partial knowledge of the ML model underlying a given NIDS and the features used to the train the model, researchers have used existing techniques such as genetic algorithm/particle swarm optimization/generative adversarial networks to generate adversarial examples [23]. An update to that work focuses on restricting the adversarial feature space so that the generated features correspond to functional network packets [24]. It does that by presenting a constrained optimization formulation for perturbing raw packet payloads while avoiding modification to the original packet function. A meta-heuristic inspired by genetic algorithm is proposed to solve the optimization problem. However, these works do not offer an explanation as to why the adversarial examples generated cause mis-classification by the underlying ML model.
Despite the presence of existing works on assessing the explainability of ML/DL and ensemble learning-based NIDSs, our work fills an important gap due to the following reasons:
  • Existing works that evaluate xAI methods on ML/DL-based NIDSs ([11, 12, 13, 14, 20, 22]): (1) do not compare the explainability of those NIDSs with that of state-of-the-art NIDSs and (2) do not include the latest xAI tools for ML/DL-based network security such as TRUSTEE [7].
  • Existing explainable-by-design AI-based attack/intrusion detection systems such as [21] are targeted at specific attacks (DNS-over-HTTPS attacks).
@@ III Background
In this paper, since we delve into explaining the decisions of state-of-the-art DL-based NIDSs, it is essential to briefly explore the types of explanation methods present in academic literature, the NIDSs that we consider for explanation in this work and the various aspects of our investigation with xAI techniques on those NIDSs.
@@ III-A Explanation Methods- Types
All explanation methods for ML/DL models can be broadly categorized into black-box and white-box explanations.
@@ III-B Prominent Black-box XAI Tools
DL-based NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the DL-based NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
Overview of TRUSTEE: TRUSTEE [7] is a state-of-the-art interpretability tool for ML/DL-based security solutions which generates interpretable decision trees (DT) for a given ML/DL model. It can be used to find out inductive biases in black-box models (e.g., shortcut learning, spurious correlations, sensitiveness to out-of-distribution samples). While generating DT, TRUSTEE focuses on the following desirable properties: model-agnosticism (should be applicable to any given black-box model), model fidelity (how accurate the DT is with respect to the original ML model), model stability (the decision rules should correctly describe how the given black-box model makes a significant number of its decisions and should be largely insensitive to the particular data samples that TRUSTEE used in the process of selecting its final DT explanation) and model complexity (decision rules should be readily recognizable by domain experts). The important aspects of generated DT explanation are summarized in a trust report which helps end users determine whether the given black-box model suffers from the problem of underspecification and can be trusted.
@@
III-C State-of-the-art Deep Learning-based NIDSs
We now present a brief overview of state-of-the-art DL-based NIDSs that we have considered in this work.
Kitsune [5] proposes an online unsupervised anomaly detection system based on an ensemble of autoencoders which is lightweight in terms of memory footprint and meant to be deployed on network gateways and routers to detect attacks on the local network. It uses a damped incremental statistics (DIS) method to extract network traffic features which has the advantage of extracting features from dynamic network traffic at high speed.
HorusEye [28] is a two-stage anomaly detection framework for IoT. In the first stage (Gulliver Tunnel), preliminary burst-level anomaly detection is implemented on the data plane using an isolation forest model converted into a set of white-listing rules. The suspicious traffic is then reported to the control plane for further investigation. To reduce the false positive rate, the control plane carries out the second stage (Magnifier), where more thorough anomaly detection is performed over the reported suspicious traffic using Kitsune DIS features and an asymmetric autoencoder with separable and dilated convolutions.
ENIDrift [29] addresses real-world issues in DL-based NIDS deployment such as concept drift (change in statistical properties of target variable over time), class-imbalanced training data and well-crafted ML attacks (training data contamination, adversarial attacks). It includes three main components: iP2V- an incremental feature extraction method based on Word2Vec, an ensemble of autoencoders, a sub-classifier generation module and an update module.
@@ III-C State-of-the-art Deep Learning-based NIDSs
The above state-of-the-art systems were selected since: (1) they were published in prominent avenues (NDSS, USENIX Security, ACSAC) and include works published as recently as 2023, (2) provide a clear performance comparison with Kitsune (which is the most cited deep learning-based NIDS with an open-source implementation) and (3) they cover a diverse ground on the application of deep learning in network intrusion detection. Kitsune uses an ensemble of autoencoders while ENIDrift builds on top of that by including a sub-classifier generation and updation module to address concept drift in real-world networks. HorusEye’s Magnifier switches the ensemble of autoencoders for a single asymmetric autoencoder to keep the solution lightweight and reduce the false-positive rate. HELAD uses a single autoencoder as well but combines it with an LSTM network.
IV xAI Analysis of Deep Learning-based NIDSs
In this work, we analyze four state-of-the-art DL-based NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible NIDS Implementation
The TRUSTEE trust report generating API has two requirements: (1) a trained ML/DL model with a predict() function taking training/test dataset as the only input and outputting predicted labels, and (2) training/test datasets and corresponding predicted labels should be in numpy array/pandas dataframe format. The kernel SHAP API shares the requirements of TRUSTEE API. Except Kitsune, other NIDS source codes do not provide trained ML model or training/test data in a format which can be directly input to TRUSTEE API or kernel SHAP API.
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML/DL models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art DL-based NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML/DL models such as the NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.